whats is the largest integer whose cube is less than 500?
Answer by oberobic(2304)

To solve inequalities, first solve the equality.
x^3 = 500
x = 500^(1/3)
Using a a calculator we find:
x = 7.93700526
.
Reviewing the question, we're looking for the largest integer whose cube is less than 500.
Obviously, 8^3 is > 500.
In fact, it is 512.
.
So, we conclude the largest integer whose cube is less than 500 must be 7.
7^3 = 343.
.
Done.